Wellplace Insurance Company processes insurance policy applications in batches of 100. One day, they had 10 batches to process and, after inspection, it was found that four batches had nonconforming policies. One batch had 3 nonconformances, another had 5, another had 2, and another had 3 nonconformance. What were (a) the proportion nonconforming for each batch, (b) the nonconformances per unit (NPU), in total for the 10 batches, and (c) the total throughput yield (TY) for the 10 batches?
Answer #1

a) Proportion of nonconforming for one batch = Number of non-conforming policies/batch size

so, Proportion of nonconforming for batches => 3/100 ,5/100,2/100,3/100 => 0.03,0.05,0.02,0.03

b) NPU per unit(NPU) in total for the 10 batches = Total no of non-conforming policies/(No of batches*batch size) = (3+5+2+3)/(10*100) = 13/1000 = 0.013

c) the total throughput yield (TY) for the 10 batches = e-NPU = e-0.013 = 0.987084135 = 98.7084135% => 98.71% (Rounding to 2 decimal places)